# Quake: Adaptive Indexing for Vector Search

### v1

@@ Abstract
Vector search, the task of finding the k-nearest neighbors of high-dimensional vectors, underpins many machine learning applications, including recommendation systems and information retrieval. However, existing approximate nearest neighbor (ANN) methods perform poorly under dynamic, skewed workloads where data distributions evolve. We introduce Quake, an adaptive indexing system that maintains low latency and high recall in such environments. Quake employs a hierarchical partitioning scheme that adjusts to updates and changing access patterns, guided by a cost model that predicts query latency based on partition sizes and access frequencies. Quake also dynamically optimizes query execution parameters to meet recall targets using a novel recall estimation model. Furthermore, Quake utilizes optimized query processing, leveraging NUMA-aware parallelism for improved memory bandwidth utilization. To evaluate Quake, we prepare a Wikipedia vector search workload and develop a workload generator to create vector search workloads with configurable access patterns. Our evaluation shows that on dynamic workloads, Quake achieves query latency reductions of 1.51.51.51.522×$\times2$ and update latency reductions of 666683×83$\times$83 × compared to state-of-the-art indexes SVS, DiskANN, HNSW, and SCANN.
@@ 1 Introduction
Vector search, the task of finding the $k$-nearest neighbors (KNN) of high-dimensional vectors, is fundamental to machine learning applications such as recommendation systems [31, 21, 42, 32, 22] and information retrieval [9, 12, 13, 35]. In these applications, each vector represents an item in a metric space, and the distance between vectors reflects semantic similarity. However, performing exact KNN search becomes computationally infeasible on large datasets due to the high dimensionality and volume of data.
To address this challenge, practitioners use approximate nearest neighbor (ANN) indexes, which trade off a controlled amount of search accuracy (recall) for significant reductions in latency. Among these indexes, there are two widely-used types: graph-based indexes and partitioned indexes.
@@
Graph-based indexes, such as HNSW [24] and DiskANN  [38, 37] construct a proximity graph where each node (vector) is connected to its approximate neighbors. Queries traverse the graph to find approximate nearest neighbors, typically achieving high recall with low latency. However, these indexes face challenges with dynamic workloads because updating the graph structure to accommodate frequent insertions and deletions is computationally intensive [44], due to the random access patterns involved in graph traversal and modification.
@@
In this work, we study the problem of minimizing query latency to meet a fixed recall target for dynamic vector search workloads, where both the queries and the base vectors can change over time. To address this problem, we develop Quake, a partitioned index system that minimizes latency by adapting the index structure to the workload. Quake’s two primary algorithmic contributions are:
First, Quake employs an adaptive hierarchical partitioning scheme that modifies the partitioning by minimizing the cost (derived from a cost model) of a query. The cost model tracks partition sizes and access frequencies as the workload is processed and determines which partitions are most negatively contributing to overall query latency. Once identified, we conduct maintenance on them by splitting or merging them based on what is expected to reduce the cost (latency). We also demonstrate our maintenance procedure is stable and converges to a local minimum of the cost model.
@@
Furthermore, Quake utilizes NUMA-aware parallelism in order to maximize memory bandwidth usage on multi-core machines.
It is a significant challenge to evaluate indexing approaches due to the lack of availability of benchmarks for online vector search. To address this challenge and comprehensively evaluate our approach, we A) prepare a Wikipedia vector search workload derived from publicly available query and update patterns of Wikipedia pages and B) develop a workload generator for creating workloads with configurable query and update patterns. We will publicly release the Wikipedia workload and workload generator as evaluation tools for the community to use. Using this, we conduct a comprehensive evaluation of Quake in comparison to seven baseline approaches.
@@
    Quake achieves the lowest search latency across all dynamic workloads in comparison to state-of-the-art graph indexes, with 1.58×1.5-8$\times$1.5 - 8 × lower search latency than HNSW and DiskANN while having 683×6-83$\times$6 - 83 × lower update latency.
@@
    We also find that APS Scanning matches the nprobe of an oracle across recall targets on Sift1M, with only a $17\%$ increase in latency relative to the oracle, and in contrast to existing methods for early termination, APS requires no offline tuning.
@@
    Quake’s NUMA-aware intra-query parallelism exhibits linear scalability and saturates memory bandwidth on the MSTuring100M dataset with 100-million vectors. Quake achieves 16×16$\times$16 × lower query latency when compared to a single-threaded version and 4×4$\times$4 × lower latency compared to a non-NUMA aware configuration.
@@ 2.1 Vector Search Workload
Recall@k is the standard metric for accuracy, defined as: $\frac{|\mathbf{G}\cap\mathbf{R}|}{k}$ where $\mathbf{R}$ is the vectors returned by the approximate search, and $\mathbf{G}$ is the ground truth set. Maintaining a consistent recall target (e.g., $>0.9$) and low latency (e.g., milliseconds per query) as both data and query patterns shift is a key challenge. The complexity of these workloads stems from their inherently dynamic and skewed nature, which few existing indexing methods handle gracefully.
@@ 2.2 Why Real-World Workloads are Hard
Table 1: Comparison of updatable vector indexes. Tuning: Requires manual parameter tuning in indexing and query processing. Maintenance: Restructures index with incremental updates. Adaptive: Utilizes query information to inform indexing.
@@ 2.3 Shortcomings of Existing Approaches
Existing indexes were often developed and evaluated under assumptions of static data distributions; conditions not met in real-world use cases. Table 1 compares a range of well-known and state-of-the-art vector indexes. Although they are widely adopted in vector databases, none fully solve the problem of maintaining low-latency, high-recall search under dynamic, skewed workloads without constant manual intervention or offline tuning.
@@
Graph-based index systems, such as HNSW [24], DiskANN [30, 37], and SVS  [5] construct a proximity graph where each node represents a vector connected to its approximate neighbors. These indexes achieve high recall with low latency in static settings by efficiently traversing the graph to locate nearest neighbors using a process known as greedy traversal. However, maintaining the graph structure under frequent insertions and deletions is computationally intensive, as each update may require rewiring multiple edges to preserve graph connectivity and proximity properties. Our evaluation (Table 1) shows that update latency can be multiple orders of magnitude higher than partitioned indexes. Additionally, graph traversal involves random memory accesses, leading to poor memory bandwidth utilization.
@@ 2.4 Technical Challenges for Partitioned Indexes
The following technical challenges have yet to be solved by existing partitioned indexes
@@
    Online Adjustment of Nprobe As the index structure and data changes, partitioned indexes need to adjust the number of partitions scanned otherwise recall will suffer. Existing early termination works are insufficient as they assume a static index and require retuning as the index and data changes.
@@ 3 Solution Overview
In Quake, vectors are partitioned into disjoint partitions (using k-means) where each partition has a representative centroid. Centroids can be further partitioned in a similar manner to add additional levels to the index. Search queries scan the index structure in a top-down fashion, finding the nearest centroids to determine the partitions to scan in the next level. Partitions in the bottom level contain the base vectors and subsets of these partitions are scanned to return the k-nearest neighbors. Utilizing a multi-level design mitigates the cost of scanning centroids, allowing us to use fine-grained partitioning of vectors at large scale, which prior work has shown to be beneficial [7].
@@
Inserts and deletes modify the Quake data structure by appending vectors to and removing vectors from index partitions. Insertions traverse the index structure top-down to find the nearest partition in the base level to the inserted vector and append to that partition. Deletes use a map to find the partition containing the vector to be deleted and the vector is removed from the partition with immediate compaction. As demonstrated in Section 2, modifications can negatively affect index performance over time, requiring maintenance (Figure 1(b)). Quake uses the following maintenance actions in order to minimize query latency:
@@
    Split Partition: Uses k-means to split a partition into two, removing the old partition and its centroid and adding two new partitions and centroids. To mitigate potential overlap due to the new partitions, we perform additional iterations of k-means over the partitions neighboring the split partitions (by centroid distance).
@@
    Add Level: Adds a level of partitioning to the index by partitioning the current top-level using k-means.
@@
Quake uses a cost model that estimates query latency to determine if maintenance actions should be taken and which partitions to apply them to. The cost model is a function of partition access patterns and sizes to determine which partitions are most contributing to the overall query latency. We check for maintenance after each operation by evaluating the cost model, but the maintenance frequency is configurable. Partitions with the largest cost contribution are considered for split or deletion. Intuitively, frequently accessed and/or large partitions are split and infrequently accessed and/or small partitions are merged as they do not justify the overhead of maintaining a centroid. See Section 4 for details on the cost model and maintenance methodology.
@@
In order to determine the number of partitions a search query should scan to reach a given recall target, we apply Adaptive Partition Scanning (APS) at each level of the index. APS solves a critical problem for partitioned indexes when applied to dynamic workloads: as the number and contents of partitions change, the number of partitions scanned needs to change, otherwise recall will degrade (Figure 1(b)). APS maintains a recall estimator based on the intermediate top-k results of the query and the geometry of neighboring partitions. As more partitions are scanned, the intermediate results and recall model are updated and when the recall estimate exceeds the target recall, the results are returned. To mitigate overheads introduced by the recall estimator, we use pre-computation of expensive geometric functions and only update the estimate when the intermediate results have changed significantly. APS supports both Euclidean and inner-product distance metrics. We cover APS in Section 5.
@@
Modern multi-core servers often use Non-Uniform Memory Access (NUMA) architectures, where memory close to a processor’s local node is faster to access than remote memory. Quake is designed to capitalize on this heterogeneous memory. It distributes index partitions across NUMA nodes. To minimize remote memory access, Quake employs affinity-based scheduling, and supports work stealing within a NUMA node to mitigate workload imbalances. By co-locating computation with the relevant data, Quake reduces remote memory accesses, saturates memory bandwidth, and thus lowers query latency. See Section 6 for further details on Quake’s NUMA-aware optimizations.
@@ 4.1 Cost Model
The cost model estimates the query latency contributed by each partition, in the index. Estimating the per-partition latency contribution enables targeted maintenance to the partitions most affecting query performance.
Partition Properties:
@@ 4.2 Conducting Maintenance
Maintenance at each level of the index proceeds in three phases—estimate, verify, and commit / reject. We first list the available actions, then derive their cost deltas, describe the workflow, and finally walk through a concrete split-verification example.
@@ 4.2.1 Maintenance Actions
After a split, refinement uses $k$-means (seeded by current centroids) on nearby partitions to mitigate overlap and ensure that each vector is assigned to its most representative partition. Nearby partitions are determined by finding the $r_{f}$ nearest centroids to the split centroids, where $r_{f}$ is a tunable parameter (typically between 10 and 100). This is a generalization of the reassignment procedure used in SpFresh [44], using additional rounds of k-means prior to reassignment, and has been applied successfully by recent index maintenance works [23] and [25]. Refinement ensures vectors are correctly assigned to their most representative partition and avoids performance degradation due to excessive overlap.
@@
If the number of centroids in the top level grows beyond a threshold, we add a new top level by clustering those centroids into fewer groups. Conversely, if the top level becomes too sparse (below a configured lower threshold), we remove the top level and merge its centroids in the level below. Both actions help maintain hierarchy balance and control centroid-scanning overhead. We defer discussion of this to our technical report.
@@ 4.2.2 Cost Deltas
We tentatively accept an action whenever $\Delta C<-\tau$. Below we give the exact $\Delta C$ formulas for the primary maintenance actions: split and merge. Full derivations are in the technical report; here we show only the final expressions and explain the terms.
Exact split delta.
@@
Exact merge delta.
@@
$\Delta\text{Split}^{\prime}_{l,j}\;=$ $\Delta O^{+}-A_{l,j}\,\lambda(s_{l,j})$ (6)
@@
and the analogous merge estimate, derived with a uniform redistribution assumption, is located in the technical report. Immediately after a tentative action we measure the actual sizes (and, for merges, the exact receiving partitions) and re-evaluate Eqs. (4) or (5). If the recomputed gain is still below $-\tau$ the action is committed; otherwise it is rolled back (§4.2.3). This “estimate-then-verify” strategy is crucial for ensuring monotonic cost improvement.
@@ 4.2.3 Decision Workflow
Maintenance is a bottom-up pass over the hierarchy. Each level executes the five stages below starting from the base level. This workflow is triggered by the user. An interesting avenue for future work is to develop scheduling and budgeting policies to call this workflow and limit its scope. In our evaluation, we trigger maintenance after a set amount of queries have been run.
Stage 0 — Track statistics.
At the end of each query batch we update, for every partition $(l,j)$: (i) size $s_{l,j}$, (ii) access count over the sliding window of queries $W$, giving $A_{l,j}=\text{hits}(l,j)/|W|$. These values feed directly into the cost formulas.
Stage 1 — Estimate.
@@
Stage 3 — Commit / Reject.
@@
Stage 4 — Propagate upward.
@@
Safety. Because every level enforces the same $\Delta<-\tau$ guard, total cost across all levels monotonically decreases and the hierarchy converges to a stable state under a fixed workload distribution (proof in technical report).

### v2

@@ Abstract
Vector search, the task of finding the k-nearest neighbors of a query vector against a database of high-dimensional vectors, underpins many machine learning applications, including retrieval-augmented generation, recommendation systems, and information retrieval. However, existing approximate nearest neighbor (ANN) methods perform poorly under dynamic and skewed workloads where data distributions evolve. We introduce Quake, an adaptive indexing system that maintains low latency and high recall in such environments. Quake employs a multi-level partitioning scheme that adjusts to updates and changing access patterns, guided by a cost model that predicts query latency based on partition sizes and access frequencies. Quake also dynamically sets query execution parameters to meet recall targets using a novel recall estimation model. Furthermore, Quake utilizes NUMA-aware intra-query parallelism for improved memory bandwidth utilization during search. To evaluate Quake, we prepare a Wikipedia vector search workload and develop a workload generator to create vector search workloads with configurable access patterns. Our evaluation shows that on dynamic workloads, Quake achieves query latency reductions of 1.51.51.51.538×38$\times$38 × and update latency reductions of 4.54.54.54.5126×126$\times$126 × compared to state-of-the-art indexes such as SVS, DiskANN, HNSW, and SCANN.
@@ 1 Introduction
Vector search, the task of finding the $k$-nearest neighbors (KNN) of a query vector against a database of high-dimensional vectors, is fundamental to modern machine learning based search [9, 11, 12, 36] and recommendation systems [32, 21, 43, 33, 22]. In these applications, a vector represents an item in a metric space, and the distance between vectors reflects semantic similarity. However, performing exact KNN search becomes computationally infeasible on large datasets due to the high dimensionality and volume of data.
To address this challenge, practitioners use approximate nearest neighbor (ANN) indexes, which trade off a controlled amount of search accuracy (recall) for significant reductions in latency. Among these, two broad classes dominate in practice: graph-based and partitioned indexes, each with distinct performance characteristics under dynamic workloads.
@@
Graph-based indexes, such as HNSW [24], DiskANN  [39, 38], and SVS  [5] construct a proximity graph where each node (vector) is connected to its approximate neighbors. Queries traverse the graph to find approximate nearest neighbors, typically achieving high recall with low latency. However, these indexes face challenges with dynamic workloads because updating the graph structure to accommodate frequent insertions and deletions is computationally intensive [45], due to the random access patterns involved in graph traversal and modification.
@@
In this work, we study the problem of minimizing query latency to meet a fixed recall target for dynamic vector search workloads, where both the queries and the base vectors can change over time. To address this problem, we develop Quake, a partitioned index for ANN search that minimizes query latency by adapting the index structure to the workload. Quake’s two primary algorithmic contributions are:
First, Quake employs an adaptive hierarchical partitioning scheme that modifies the partitioning by minimizing the cost (derived from a proposed cost model) of a query. The cost model tracks partition sizes and access frequencies as the workload is processed and determines which partitions are most negatively contributing to overall query latency. Once identified, we split or merge these partitions based on expected cost reduction derived from our proposed cost model. We also demonstrate that our maintenance procedure is stable and converges to a local minimum of the cost model.
@@
Furthermore, Quake utilizes NUMA-aware parallelism to maximize memory bandwidth usage on multi-core machines.
It is a significant challenge to evaluate indexing approaches due to the lack of availability of benchmarks for online vector search. To address this challenge and comprehensively evaluate our approach, we A) prepare a Wikipedia vector search workload derived from publicly available query and update patterns of Wikipedia pages and B) develop a workload generator for creating workloads with configurable query and update patterns. We will publicly release the Wikipedia workload and workload generator as evaluation tools for the community to use. Using these, we conduct a comprehensive evaluation of Quake in comparison to seven baseline approaches.
@@
    Quake achieves the lowest search time across all dynamic workloads compared to state-of-the-art graph indexes, with $1.5$-13×13$\times$13 × lower search latency than HNSW, DiskANN, and SVS while having $18$-126×126$\times$126 × lower update latency.
@@
    We also find that APS matches the nprobe of an oracle across recall targets on Sift1M, with only a $17$-$29\%$ increase in latency relative to the oracle.
@@
    APS performs on-par or better than existing early termination methods [48, 7, 18] and requires no offline tuning.
  4. 4.
    Quake’s NUMA-aware query processing exhibits linear scalability and high memory bandwidth utilization on the MSTuring100M dataset. Quake achieves 20×20$\times$20 × and 4×4$\times$4 × lower query latency compared to single-threaded and non-NUMA aware configurations, respectively.
@@ 2.1 Vector Search Workload
Recall@k is the standard metric for accuracy, defined as: $\frac{|\mathbf{G}\cap\mathbf{R}|}{k}$ where $\mathbf{R}$ is the vectors returned by the approximate search, and $\mathbf{G}$ is the ground truth set. Maintaining a consistent recall target (e.g., $>90\%$) and low latency (e.g., milliseconds per query) as both data and query patterns shift is a key challenge. The complexity of these workloads stems from their inherently dynamic and skewed nature, which few existing indexing methods handle gracefully.
@@ 2.2 Why Real-World Workloads are Hard
Table 1: Comparison of updatable vector indexes. Tuning: Requires manual parameter tuning in indexing/query processing. Maintenance: Modifies index with incremental updates. Adaptive: Uses query information to inform indexing.
@@ 2.3 Shortcomings of Existing Approaches
Existing indexes were often developed and evaluated under assumptions of static data distributions; conditions not met in real-world use cases. Table 1 compares a range of state-of-the-art vector indexes. Although widely adopted in vector databases, none fully solve the problem of maintaining low-latency, high-recall search under dynamic, skewed workloads without constant manual intervention or offline tuning.
@@
Graph-based index systems, such as HNSW [24], DiskANN [31, 38], and SVS  [5] construct a proximity graph where each node represents a vector connected to its approximate neighbors. These indexes achieve high recall with low latency in static settings by efficiently traversing the graph to locate nearest neighbors using a process known as greedy traversal. However, maintaining the graph structure under frequent updates is computationally intensive, as each update may require rewiring multiple edges to preserve graph connectivity and proximity properties. Our evaluation (Table 1) shows that update latency can be multiple orders of magnitude higher than partitioned indexes.
@@ 2.4 Technical Challenges for Partitioned Indexes
The following technical challenges are yet to be solved by existing partitioned indexes
@@
    Online Adjustment of Nprobe As the index structure and data change, partitioned indexes need to adjust the number of partitions scanned or recall will suffer. Existing early termination works are insufficient as they assume a static index and require retuning as the index and data change.
@@ 3 Solution Overview
Quake organizes the vectors in a multi-level index, where each level is a partitioned index similar to Faiss-IVF [8]. The lowest level in the index is constructed by organizing the vectors into disjoint partitions (using k-means clustering) where each partition has a representative centroid. These centroids can be further partitioned in a similar manner to create additional levels in the index. Search queries scan the index in a top-down fashion, finding the nearest centroids at each level to determine the partitions to scan in the next level. Partitions in the lowest level contain the actual vectors and subsets of these partitions are scanned to return the k-nearest neighbors. Utilizing a multi-level design enables us to employ fine-grained partitioning of vectors at large scale (shown to improve search quality [7]), while mitigating the high cost of scanning centroids.
@@
Inserts and deletes modify the Quake data structure by appending vectors to and removing vectors from index partitions. Insertions traverse the index structure top-down to find the nearest partition in the lowest level to the inserted vector and append to that partition. Deletes use a map to find the partition containing the vector to be deleted and the vector is removed from the partition with immediate compaction. As demonstrated in Section 2, modifications can negatively affect index performance over time, requiring maintenance (Figure 1(b)). Quake uses the following maintenance actions in order to minimize query latency:
@@
    Split Partition: Uses k-means clustering to split a partition into two, removing the old partition and its centroid and adding two new partitions and centroids. To mitigate potential overlap due to the new partitions, we perform additional iterations of k-means clustering over the partitions neighboring the split partitions (by centroid distance).
@@
    Add Level: Adds a level of partitioning to the index by partitioning the current top-level using k-means clustering.
@@
Quake uses a cost model that estimates query latency to determine if maintenance actions should be taken and which partitions to apply them to. The cost model is a function of partition access patterns and sizes to determine which partitions are contributing most to the overall query latency. We check for maintenance after each operation by evaluating the cost model, but the maintenance frequency is configurable. Partitions with the largest cost contribution are considered for split or deletion. Intuitively, frequently accessed and/or large partitions are split and infrequently accessed and/or small partitions are merged as they do not justify the overhead of maintaining a centroid. See Section 4 for details on the cost model and maintenance methodology.
@@
In order to determine the number of partitions a search query should scan to reach a given recall target, we apply Adaptive Partition Scanning (APS) at each level of the index. APS solves a critical problem for partitioned indexes when applied to dynamic workloads: as the number and contents of partitions change, the number of partitions scanned needs to change, otherwise recall will degrade (Figure 1(b)). APS maintains a recall estimator based on the intermediate top-k results of the query and the geometry of neighboring partitions. As more partitions are scanned, the intermediate results and recall model are updated and when the recall estimate exceeds the target recall, the results are returned. To mitigate overheads introduced by the estimator, we use pre-computation of expensive geometric functions and only update the estimate when the intermediate results have changed significantly. APS supports both Euclidean and inner-product distance metrics. We cover APS in Section 5.
@@
Modern multi-core servers often use Non-Uniform Memory Access (NUMA) architectures, where memory close to a processor’s local node is faster to access than remote memory. Quake is designed to capitalize on this heterogeneous memory. It distributes index partitions across NUMA nodes. To minimize remote memory access, Quake employs affinity-based scheduling, and supports work stealing within a NUMA node to mitigate workload imbalances. By co-locating computation with the relevant data, Quake reduces remote memory accesses, saturates memory bandwidth, and thus lowers query latency. See Section 6 for details on Quake’s NUMA-aware optimizations.
@@ 4.1 Cost Model
The cost model estimates the query latency contributed by each partition in the index. Estimating the per-partition latency contribution enables targeted maintenance to the partitions most affecting query performance.
Partition Properties
@@ 4.2 Conducting Maintenance
Maintenance at each level of the index proceeds in three phases—estimate, verify, and commit / reject. We first list the available actions, then derive their cost deltas, describe the workflow, and finally walk through a concrete example.
@@ 4.2.1 Maintenance Actions
After a split, refinement uses $k$-means (seeded by current centroids) on nearby partitions to mitigate overlap and ensure that each vector is assigned to its most representative partition. Nearby partitions are determined by finding the $r_{f}$ nearest centroids to the split centroids, where $r_{f}$ is a tunable parameter (typically between 10 and 100). This is a generalization of the reassignment procedure used in SpFresh/LIRE [45], using additional rounds of k-means prior to reassignment, and has been applied successfully by recent index maintenance works [23] and [25]. Refinement avoids performance degradation by mitigating overlap and ensuring vectors are assigned to their most representative partition.
@@
If the number of centroids in the top level grows beyond a threshold, we add a new top level by clustering those centroids. Conversely, if the top level becomes too sparse (below a configured lower threshold), we remove the top level and merge its centroids in the level below. Both actions help maintain hierarchy balance and control centroid-scanning overhead. We defer discussion of this to our technical report [27].
@@ 4.2.2 Cost Deltas
We tentatively accept an action whenever $\Delta C<-\tau$. Below we give the exact $\Delta C$ formulas for the primary maintenance actions: split and merge. Full derivations are in the technical report [27]; here we show only the final expressions.
Exact Split Delta
@@
Exact Merge Delta
@@
$\Delta^{\prime}\text{Split}_{l,j}\;=$ $\Delta O^{+}-A_{l,j}\,\lambda(s_{l,j})$ (6)
@@
and the analogous merge estimate, derived with a uniform redistribution assumption, is located in the technical report [27]. Immediately after a tentative action we measure the actual sizes (and, for merges, the exact receiving partitions) and re-evaluate Eqs. (4) or (5). If the recomputed gain is still below $-\tau$ the action is committed; otherwise it is rolled back (§4.2.3). This “estimate-then-verify” strategy is crucial for ensuring monotonic cost improvement.
@@ 4.2.3 Decision Workflow
Maintenance is a bottom-up pass over the hierarchy. Each level executes the five stages below starting from the base level. This workflow is triggered by the user. An avenue for future work is to develop scheduling policies to call this workflow and limit its scope. In our evaluation, we trigger maintenance after a set amount of queries have been run.
Stage 0: Track Statistics
At the end of each query batch we update, for every partition $(l,j)$: (i) size $s_{l,j}$, (ii) access count over the sliding window of queries $W$, giving $A_{l,j}=\text{hits}(l,j)/|W|$. These values are inputs of the cost model.
Stage 1: Estimate
@@
Stage 3: Commit / Reject
@@
Stage 4: Propagate Upward.
@@
Safety: Because every level enforces the same $\Delta<-\tau$ guard, total cost across all levels monotonically decreases and the hierarchy converges to a stable state under a fixed workload distribution (proof in technical report [27]).